Calculate the percentage by mass of the element adipic acid, C6 H10 O4
1
Expert's answer
2013-05-17T10:36:40-0400
The percentage we can calculate by formula:

W(E) = n*Ar(E)/Mr(substance)
where
W(E) - the percentage of element
n - amount of element in substance
Ar(E) - atomic weight of element
Mr(substance) - molecular weight of substance
Mr(substance) = 12*6+1*10+16*4 = 146 amu
W(C) = 6 * 12/ 146 = 0.493 * 100 % = 49.3%
W(H) = 1 * 10/ 146 = 0.0684 * 100 % = 6.84%
W(O) = 16 * 4/ 146 = 0.438 * 100 % = 43.8%
Answer: W(C) = 49.3%, W(H) = 6.84%, W(O) = 43.8%